# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v7

@@ Abstract
Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph. In this paper, we first present a novel data structure, GLOVER (Graph Level Order), which partitions nodes into a total ordering of containers. This data structure establishes a well-ordering on oriented graphs and allows for the construction of a decreasing sequence of subsets of nodes. This sequence proves the non-existence of counterexamples to the SSNC and precisely identifies the required node. Further, our approach finds the occurrence of dense regular graphs inside containers. This finding extends the SSNC to the discovery of multiple nodes satisfying the degree doubling property. Beyond theoretical implications, the algorithm and data structure have practical applications in network optimization and algorithm design.
@@ 1 Introduction
Despite decades of research, the SSNC remains one of graph theory’s challenging open problems. Special cases of the conjecture have been explored extensively. Kaneko and Locke [11] showed that the SSNC holds for graphs with a minimum degree of at most six. Brantner et al. [1] demonstrated that if the conjecture were false, infinitely many counterexamples would exist. Chen et al. [3] established a lower bound, proving that every oriented graph contains a vertex whose second neighborhood is at least $lambda$ times larger than its first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1. In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture. This was an extension of the work done by Botler et al. [2], where their work confirms that almost all orientations of $G(n,p)$ satisfy the SSNC.
@@
This paper builds on these foundations, generalizing an approach initially developed for the minimum degree 7 case [9]. Our key contribution is the introduction of a novel data structure, GLOVER (Graph Level Order), which partitions nodes into a total ordering of containers. This structure enables the construction of a decreasing sequence of subsets of nodes, forming the basis of our algorithmic proof of the SSNC. Our approach not only proves the conjecture but also identifies the specific node(s) whose out-degree doubles in the graph’s square.
Beyond its theoretical significance, the SSNC has practical applications in domains like social networks, epidemiology, and algorithm design. For instance, in the context of disease control, the conjecture helps identify individuals whose influence in their second-degree connections surpasses their direct contacts, offering insights into the spread of infections.
Unlike traditional proof-theoretic approaches, this work emphasizes the development of an efficient algorithm to locate the target node. We also analyze the complexity of our algorithm, demonstrating its practical feasibility for large-scale problems. The paper includes extensive supporting materials, such as detailed proofs, examples, and an interactive website to visualize key ideas, ensuring accessibility and transparency.
The remainder of this paper is structured as follows: Section 2 introduces common terminology used in graph theory. Section 3 outlines our approach to the problem, while Section 4 provides illustrative examples. Section 5 focuses on conditions under which a node’s out-degree may fail to double, and Section 6 examines the role of back arcs in oriented graphs. Section 7 explores the conditions that lead to doubling of interior degrees, examining their significance in relation to the SSNC. It introduces a divide-and-conquer approach, where the graph is partitioned, and each subproblem is solved independently before being integrated into the broader structure. Section 8 explores density constraints in containers. Section 9 investigates the impact of transitive triangles and Seymour diamonds on our algorithm. Finally, Section 10 presents the full Decreasing Sequence Property algorithm and its proof of the conjecture.
@@ 2 Graph Theory Terminology
(Square version of Seymour’s Second Neighborhood Conjecture). Prove that for every oriented graph, $G$, there exists a vertex whose out-degree at least doubles when you square the oriented graph.
Now that we have two different conjectures, we can investigate how they relate to one another. Conjecture 1.1 relates the notion of a node’s degree doubling to its neighborhoods. Similarly, conjecture 2.1 relates this same concept to the square graph. The defining characteristics of these conjectures are the relationship between a node’s out-degree at least doubling, and its second neighborhood being greater than or equal to the first. By the definition of an arc in the square graph, we can see that an arc is defined in this square graph only for paths of length two in the original graph. This is why these two conjectures are seen as identical.
@@
Next, we turn our attention to the term degree-doubling nodes. By this term, we aim to get to the heart of conjecture 1.1. That conjecture is looking for a node whose degree at least doubles. In order to do this, we examine the node’s first and second neighborhoods. Then identify a node whose second neighborhood is at least as large as the first, as required by the conjecture. A similar definition, a Seymour vertex, fulfills conjecture 2.1 - it’s out-degree in the squared graph is at least double its original out-degree. We could equivalently search for a Seymour vertex or a degree-doubling node.
@@ 3 Approach and Motivation
We begin our analysis by considering general oriented graphs. We make no initial assumptions about node degrees or graph properties. There are, however, fundamental characteristics common to all graphs. We know that there is an out-degree distribution. Consequently, every oriented graph will have a minimum out-degree $d$, representing the lowest out-degree among all nodes. By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
@@
When the graph is partitioned, even larger graphs, often with higher value minimum degree nodes and thus more partitions, are able to be broken into subproblems. These subproblems are then able to be solved independent of one another. These subproblems are each their own instances of the SSNC, just a localized version. What the partition is doing is breaking a node’s individual degree into two parts—the degree within the container and the degree to the next container. The degree within the container is the subproblem that will be solved separately, and the degree to the next container is how they connect back to the global problem.
@@ 4 Initial Lemmas
Minimum Out-Degree $<$ 3 Suppose that a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3. Then $v_{0}$’s out-degree will at least double.
@@
Case 3: Assume that $d^{+}(v_{0})=2$. Then $v_{0}$ is connected to two nodes, $v_{1}$ and $v_{2}\in G$. We know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$. Consider the induced subgraph $G[v_{1},v_{2}]$. There is at most one directed edge between $v_{1}$ and $v_{2}$ in this induced subgraph. This means that one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes. Let these nodes be $v_{3}$ and $v_{4}$. The node $v_{0}$ cannot be one of these nodes. Hence $v_{0}$ will have at least two second neighbors - $v_{3}$ and $v_{4}$, which makes its second neighborhood at least as large as its first. Thus, its out-degree will double in $G^{2}$.
Therefore, in all cases, $v_{0}$’s out-degree at least doubles.
@@
$\Rightarrow$ 111122223333 $\Rightarrow$ 45678 $\Rightarrow$ H
@@
5 Theoretical Foundation
We will introduce some new terms to help with the rest of this paper. The partitions we discussed earlier will now be formally referred to as containers. Additionally, we shall discuss the decreasing sequence property and its significance. After that, we will discuss interior, exterior, and back arcs. We will show the importance of these arcs with some foundational lemmas. The GLOVER data structure will then be used to wrap up.
@@ 5 Theoretical Foundation
Figure 4: Illustration of containers, each enclosed in a checkered box. Containers group nodes based on their distance from the minimum degree node, with no assumptions on container relationships or sizes.
@@
$V(G)$ denotes the graph $G$’s vertex set and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
@@
We will investigate the structure of oriented graphs from two angles by using the idea of containers. Alternatively, we will investigate it from the viewpoint of containers or the minimum degree node. Containers are seen based on their distance from the minimum degree node. Nodes are seen based on their position relative to containers. They are either before a container, in a container, or after a container.
Specifically, the case where $d^{+}(v_{0})=3$ and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) are important because they provide information about what a counterexample to the conjecture would involve. A graph $G$ was constructed in which the out-degree of $v_{0}$ does not double in $G^{2}$. A simple graph of degree three corresponds to the smallest cycle that can exist in an undirected graph. Therefore, if the edges are oriented in the same direction, a cycle of order three can be created among the nodes adjacent to $v_{0}$. If we look at this situation further, we see that in this example, a node that is an out-neighbor of $v_{0}$ had its out-degree (at least) double. This section will initiate the examination of the adequate conditions for other nodes in the graph to not have their degrees double.
However, we will need to define new concepts to proceed with this line of reasoning. To prevent a node’s out-degree from doubling, we must first define a term. The decreasing sequence property is attributed to a node in $G$ whose out-degree does not double in $G^{2}$. There are instances where $v_{0}$ lacks its out-degree double in $G^{2}$, as we have seen. We’ll investigate which other nodes can possess the decreasing sequence property and which network features are necessary for this to occur.
Example 5.2.
0
$\Rightarrow$ 1111222233334444d $\Rightarrow$ d + $1$d + $2$d + 33332*d-1 $\Rightarrow$ H
@@
If the size of a node $x\in V(G)$ in an oriented graph $G$ is strictly bigger than the size of its second out-neighborhood, that is, $|N^{+}(x)|>|N^{++}(x)|$, then $x$ is said to have the decreasing sequence property.
Two competing new ideas have been presented, the decreasing sequence property (DSP) and the minimal degree node. A crucial starting point for our approach and degree values is a minimal degree node. It acts as the anchor by being chosen as the beginning point. The remaining nodes are currently being forced into the attributes that will surface in the counterexample construction by the minimal degree node. It acts as a catalyst, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
In contrast, the DSP introduces constraints on connectivity. This concept is akin to a bottleneck. A bottleneck in graph theory represents a point that limits flow or connectivity within the network, a role precisely filled by the DSP. Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively.
Lemma 5.1.
Total Order Induced by Containers Relative to the Minimum Degree Node
Let $G$ be a graph with a minimum-degree node $v_{0}$. Construct a series of containers $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{k}$ ordered by distance relative to $v_{0}$. Each container $C_{i}$ is a set of nodes at distance $i$ from $v_{0}$. This imposes a total order on the nodes in $G$. This means that every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
Proof.
To prove a total order, we need to show that any two nodes in $G$ can be compared. Let $u,v$ be two nodes in $G$ such that $u\neq v$. We assume that both $u$ and $v$ are reachable from $v_{0}$. Distance is a total ordering, so one of three things is true: either $dist(v_{0},u)<dist(v_{0},v)$, $dist(v_{0},u)=dist(v_{0},v)$, or $dist(v_{0},u)>dist(v_{0},v)$. The situation where $dist(v_{0},u)<dist(v_{0},v)$ implies that $u$ is in a closer container than $v$. The situation where $dist(v_{0},u)=dist(v_{0},v)$ implies that the nodes $u\text{ and }v$ are in the same container. And the situation $dist(v_{0},u)>dist(v_{0},v)$ implies that $u$ is in a further container. This shows that it is a total order. ∎
While imposing order on our graph is a solid first step, the containers do more than just enforce overall order. In essence, they allow for the grouping of comparisons between nodes that are equally spaced from a reference node, $v_{0}$. A typical total order, like lexicographical order, does not take this into consideration. This ”containerized ordering” allows us to differentiate nodes based on particular adjacency qualities inside each distance group in addition to their distance from $v_{0}$.
Definition 5.3.
A Level Total Order (LTO) on a set $S$ can be defined as follows:
  • Total Order: For any two elements $u$ and $v$ within the set $S$, exactly one of the following must hold:
    • $u$ is less than $v$ (denoted as $u<v$),
    • $u$ is greater than $v$ (denoted as $u>v$), or
    • $u$ is equal to $v$ (denoted as $u=v$).
  • Leveled Structure:
    • The elements of $S$ are partitioned into a leveling of containers. $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{n}$
    • Each container itself forms a subset of $S$.
    • The containers are sequentially ordered relative to each other.
  • Comparability within Containers: For any two elements $u,v\in C_{i}$, their order can be determined based on a specific criterion.
This order means that we have order in two ways. First, the containers themselves are ordered. Second, the nodes within the containers are ordered. This ordering must also be consistent within the containers.
Lemma 5.2.
Definition 5.4.
Let $x$ be a node with the decreasing sequence property and let $y\in N^{+}(x)$. The interior neighbors of $y$ with respect to $x$, denoted $int(x,y)$, are the common out-neighbors of both $x$ and $y$:
@@ A Level Total Order (LTO) on a set $S$ can be defined as follows:
The number of interior neighbors is the interior degree of $y$ with respect to $x$.
@@
Let $x$ be a node with the decreasing sequence property, and let $y\in N^{+}(x)$. The exterior neighbors of $y$ with respect to $x$, denoted $ext(x,y)$, are the out-neighbors of $y$ that are second neighbors of $x$:
@@
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. To do so, though, we need to consider the option of back arcs and see how they might hurt our algorithm. These have the possibility of altering a container’s distance to the minimum degree node. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes.
Having defined our arcs, we can talk about the decreasing sequence property. This property is explaining the nature of how the exterior degrees, and thus the sizes of the containers, decrease as we get further from the minimum degree node. Exterior degrees, though, are a part of a node’s degree equation. There’s another part of it though. As we move further from the minimum degree node, the size of the interior degrees increases.
This phenomenon reflects a yin and yang relationship between the interior and exterior degrees. In all of this, the minimum degree node is serving as a catalyst and driving force for the process. At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s out-degrees. With this definition of interior and exterior nodes now introduced, we are now ready to introduce the Graph Level Order (GLOVER) data structure.
@@ A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
To establish a graph level order, we first need a minimum degree node. So we take the minimum value on the degree vector of nodes. We will call that node $v_{0}$. Knowing the minimum degree node of this graph, we can partition the nodes of this graph into containers, $\{C_{1},C_{2},\dots,C_{k}\}$. Each of these containers $C_{i}=\{u|d(v_{0},u)=i\}$ is a set of nodes at distance $i$ from $v_{0}$. These containers can then be sorted based on the distance $i$ into a total order for the GLOVER requirement.
@@
  • Leveled Structure with Containers - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ each correspond to nodes at a specific distance from $v_{0}$.
  • Comparability within Containers - For any two nodes $u$ and $v$ in the same container $C_{i}$, we can select a second metric to distinguish $u$ from $v$.
@@
What the GLOVER data structure does is transform the partitions and distance metric into an ordering for containers. This ordering is more than just representative, though. The containers help to organize the nodes. What this suggests is that each partition, or set of nodes, is being totally ordered as we move away from the chosen minimum degree node. This is a clear distinction from standard graph theory. Graph nodes are generally thought of as able to be placed anywhere in 2-dimensional space. However, in order for this data structure and algorithm to work, nodes need to be placed in proper alignment with their distance from the chosen minimum degree node.
@@
GLOVER Supports Induction In any oriented graph, without back arcs, the minimum degree node $v_{0}$ within a GLOVER structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
@@
We have already shown that the containers in GLOVER form a total order. Further, they are a discrete total order. Because of this, they are a well ordering. A set is considered well-ordered if every non-empty subset of that set has a least element. A discrete total order is a total order where each element has a distinct ”next” element in the order. Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a ”smallest” element according to the order. This fulfills the condition of being well-ordered. Every non-empty subset will have a least element, which is the defining characteristic of a well-ordered set. ∎
This becomes an important concept because now, in addition to the nodes, we also have the containers as subsets of $S$. These containers are totally ordered and thus can also be used as a means of mathematical induction. We now have several options of how to proceed with our proofs, but the foundation will be the Graph Level Order (GLOVER) data structure, which orders the nodes based on their distance from the chosen minimum degree node $v_{0}$.

### v8

@@ Abstract
Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph.
In this paper, we first present a novel data structure, GLOVER (Graph Level Order), which partitions nodes into a total ordering of containers.
This data structure establishes a well-ordering on oriented graphs and allows for the construction of a decreasing sequence of subsets of nodes.
This sequence proves the non-existence of counterexamples to the SSNC and precisely identifies the required node.
Further, our approach finds the occurrence of dense regular graphs inside containers.
This finding extends the SSNC to the discovery of multiple nodes satisfying the degree doubling property.
Beyond theoretical implications, the algorithm and data structure have practical applications in network optimization and algorithm design.
@@ 1 Introduction
Despite decades of research, the SSNC remains one of graph theory’s challenging open problems.
Special cases of the conjecture have been explored extensively.
Kaneko and Locke [11] showed that the SSNC holds for graphs with a minimum degree of at most six.
Brantner et al. [1] demonstrated that if the conjecture were false, infinitely many counterexamples would exist.
Chen et al. [3] established a lower bound, proving that every oriented graph contains a vertex whose second neighborhood is at least $lambda$ times larger than its first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1.
In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture.
This was an extension of the work done by Botler et al. [2], where their work confirms that almost all orientations of $G(n,p)$ satisfy the SSNC.
@@
This paper builds on these foundations, generalizing an approach initially developed for the minimum degree 7 case [9].
Our key contribution is the introduction of a novel data structure, GLOVER (Graph Level Order), which partitions nodes into a total ordering of containers.
This structure enables the construction of a decreasing sequence of subsets of nodes, forming the basis of our algorithmic proof of the SSNC.
Our approach not only proves the conjecture but also identifies the specific node(s) whose out-degree doubles in the graph’s square.
Beyond its theoretical significance, the SSNC has practical applications in domains like social networks, epidemiology, and algorithm design.
For instance, in the context of disease control, the conjecture helps identify individuals whose influence in their second-degree connections surpasses their direct contacts, offering insights into the spread of infections.
Unlike traditional proof-theoretic approaches, this work emphasizes the development of an efficient algorithm to locate the target node.
We also analyze the complexity of our algorithm, demonstrating its practical feasibility for large-scale problems.
The paper includes extensive supporting materials, such as detailed proofs, examples, and an interactive website to visualize key ideas, ensuring accessibility and transparency.
The remainder of this paper is structured as follows: Section 2 introduces common terminology used in graph theory. Section 3 outlines our approach to the problem, while Section 4 provides illustrative examples. Section 5 introduces a data structure that will be instrumental for the proof of the conjecture. Section 6 begins to use that data structure to establish some foundational lemmas and properties. Section 7 examines the role of back arcs in oriented graphs. Section 8 explores the conditions that lead to doubling of interior degrees, examining their significance in relation to the SSNC. It introduces a divide-and-conquer approach, where the graph is partitioned, and each subproblem is solved independently before being integrated into the broader structure. Section 9 explores density constraints in containers. Section 10 investigates the impact of transitive triangles and Seymour diamonds on our algorithm. Finally, Section 11 presents the full Decreasing Sequence Property algorithm and its proof of the conjecture.
@@ 2 Graph Theory Terminology
(Square version of Seymour’s Second Neighborhood Conjecture).
Prove that for every oriented graph, $G$, there exists a vertex whose out-degree at least doubles when you square the oriented graph.
Now that we have two different conjectures, we can investigate how they relate to one another. Conjecture 1.1 relates the notion of a node’s degree doubling to its neighborhoods. Similarly, conjecture 2.1 relates this same concept to the square graph. The defining characteristics of these conjectures are the relationship between a node’s out-degree at least doubling and its second neighborhood being greater than or equal to the first. By the definition of an arc in the square graph, we can see that an arc is defined in this square graph only for paths of length two in the original graph. This is why these two conjectures are seen as identical.
@@
Next, we turn our attention to the term degree-doubling nodes. By this term, we aim to get to the heart of conjecture 1.1. That conjecture is looking for a node whose degree at least doubles. In order to do this, we examine the node’s first and second neighborhoods. Then identify a node whose second neighborhood is at least as large as the first, as required by the conjecture. A similar definition, a Seymour vertex, fulfills conjecture 2.1—its out-degree in the squared graph is at least double its original out-degree. We could equivalently search for a Seymour vertex or a degree-doubling node.
@@ 3 Approach and Motivation
We begin our analysis by considering generally oriented graphs. We make no initial assumptions about node degrees or graph properties. There are, however, fundamental characteristics common to all graphs. We know that there is an out-degree distribution. Consequently, every oriented graph will have a minimum out-degree $d$, representing the lowest out-degree among all nodes. By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
@@
When the graph is partitioned, even larger graphs, often with higher value minimum degree nodes and thus more partitions, are able to be broken into subproblems. These subproblems are then able to be solved independently of one another. These subproblems are each their own instances of the SSNC, just a localized version. What the partition is doing is breaking a node’s individual degree into two parts—the degree within the container and the degree to the next container. The degree within the container is the subproblem that will be solved separately, and the degree to the next container is how they connect back to the global problem.
@@ 4 Initial Lemmas
Minimum Out-Degree $<$ 3
Suppose that a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3. Then $v_{0}$’s out-degree will at least double.
@@
Case 3: Assume that $d^{+}(v_{0})=2$. Then $v_{0}$ is connected to two nodes, $v_{1}$ and $v_{2}\in G$. We know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$. Consider the induced subgraph $G[v_{1},v_{2}]$. There is at most one directed edge between $v_{1}$ and $v_{2}$ in this induced subgraph. This means that one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes. Let these nodes be $v_{3}$ and $v_{4}$. The node $v_{0}$ cannot be one of these nodes. Hence $v_{0}$ will have at least two second neighbors—$v_{3}$ and $v_{4}$, which makes its second neighborhood at least as large as its first. Thus, its out-degree will double in $G^{2}$.
Therefore, in all cases, $v_{0}$’s out-degree at least doubles. ∎
@@
$\Rightarrow$ 111122223333 $\Rightarrow$ 45678 $\Rightarrow$
@@
5 Building a Data Structure
In this section, we develop a new data structure for oriented graphs (Graph Level Order). This structure is specifically designed for the SSNC and oriented graphs but will have applications beyond. In order to define this data structure, we will need to formally refer to the partitions we discussed earlier as containers. Additionally, we shall discuss interior, exterior, and back arcs and show the influence they have on our data structure. Finally, we will speak about how this data structure can use the divide and conquer paradigm to help solve the SSNC.
@@ 5 Building a Data Structure
Figure 4: Illustration of containers, each enclosed in a checkered box. Containers group nodes based on their distance from $v_{0}$, with no assumptions on container relationships or sizes.
@@
$V(G)$ denotes the graph $G$’s vertex set, and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
@@
We will investigate the structure of oriented graphs from two angles—from the viewpoint of containers or $v_{0}$. Containers are analyzed based on their distance from $v_{0}$. Nodes are seen based on their position within these containers. They are either before a container, in a container, or after a container.
@@
Let $x$ be a node in $G$ and let $y\in N^{+}(x)$. The interior neighbors of $y$ with respect to $x$, denoted $int(x,y)$, are the common out-neighbors of both $x$ and $y$:
@@
The number of interior neighbors is the interior degree of
$y$ with respect to $x$.
@@
Let $x$ be a node in $G$, and let $y\in N^{+}(x)$. The exterior neighbors of $y$ with respect to $x$, denoted $ext(x,y)$, are the out-neighbors of $y$ that are second neighbors of $x$:
@@
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. To do so, though, we need to consider the option of back arcs and see how they might hurt our algorithm. These have the possibility of altering a container’s distance to $v_{0}$. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes. We will speak more about back arcs in the next section.
@@ A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
To establish a graph level order, we first need a minimum degree node. So we take the minimum value on the degree vector of nodes. We will call that node $v_{0}$.
Knowing the minimum degree node of this graph, we can partition the nodes of this graph into containers, $\{C_{1},C_{2},\dots,C_{k}\}$.
Each of these containers $C_{i}=\{u|d(v_{0},u)=i\}$ is a set of nodes at distance $i$ from $v_{0}$. These containers can then be sorted based on the distance $i$ into a total order for the GLOVER requirement.
@@
  • Leveled Structure with Containers—By construction, the containers $C_{1},C_{2},\dots,C_{k}$ each correspond to nodes at a specific distance from $v_{0}$.
  • Comparability within Containers—For any two nodes $u$ and $v$ in the same container $C_{i}$, we can select a second metric to distinguish $u$ from $v$.
@@
What the GLOVER data structure does is transform the containers and distance metric into an ordering for containers. This ordering is more than just representative, though. The containers help to organize the nodes. What this suggests is that each partition, or set of nodes, is being totally ordered as we move away from $v_{0}$. This is a clear distinction from standard graph theory. Graph nodes are generally thought of as able to be placed anywhere in 2-dimensional space. However, in order for this data structure and algorithm to work, nodes need to be placed in proper alignment with their distance from $v_{0}$.
@@
GLOVER Supports Induction
In any oriented graph, without back arcs, the minimum degree node $v_{0}$ within a GLOVER structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
@@
We have already shown that the containers in GLOVER form a total order. Further, they are a discrete total order. Because of this, they are a well ordering.
A set is considered well-ordered if every non-empty subset of that set has a least element.
A discrete total order is a total order where each element has a distinct ”next” element in the order.
Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a ”smallest” element according to the order. This fulfills the condition of being well-ordered.
Every non-empty subset will have a least element, which is the defining characteristic of a well-ordered set.
This becomes an important concept because now, in addition to the nodes, we also have the containers as subsets of $S$. These containers are totally ordered and thus can also be used as a means of mathematical induction.
We now have several options of how to proceed with our proofs, but the foundation will be the Graph Level Order (GLOVER) data structure, which orders the nodes based on their distance from $v_{0}$.
One of the additional benefits of partitioning the oriented graph into containers with interior arcs is that each container resembles its own local SSNC. In this instance, the interior arc has an interior degree, and if we are able to find a positive solution to this instance, we can return to the global problem with a node whose interior degree has doubled. Then, under the assumption of no back arcs, we would only need to prove that this node has its exterior degree double to prove that it is a solution to the SSNC. This is another benefit of the GLOVER data structure, as it provides us with a means to execute divide-and-conquer.
This does not only help with the search for the degree-doubling node. As we shall see in the coming sections, there are multiple lemmas that will be dependent on the condition of an interior degree doubling. What that is stating is that the local subproblem has a solution to the SSNC. This overall proof strategy is through contradiction. We are attempting to prove the SSNC by assuming that it is not true and observing the patterns that must be present in the resulting oriented graphs. As we observe these patterns that must be present, we will continue to keep a log until we arrive at a contradiction.
6 Using the Data Structure
Having defined the GLOVER data structure, we are ready to begin using it. The first and most important concept that we will define that will guide our research is the decreasing sequence property. With that and the GLOVER data structure, and the ammunition of knowing that any oriented graph without back-arcs can be represented in a GLOVER, we can assume we are working in this environment.
The decreasing sequence property explains the nature of how the exterior degrees, and thus the sizes of the containers, decrease as we get further from $v_{0}$. Exterior degrees, though, are a part of a node’s degree equation. There’s another part of it, though. As we move further from $v_{0}$, if there are no back arcs, the size of the interior degrees must increase.
This phenomenon reflects a yin and yang relationship between the interior and exterior degrees. In all of this, the minimum degree node, $v_{0}$, is serving as a catalyst and driving force for the process. At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s out-degrees.
The case where $d^{+}(v_{0})=3$ and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) are important because they provide information about what a counterexample to the conjecture would involve. A graph $G$ was constructed in which the out-degree of $v_{0}$ does not double in the square $G^{2}$. A simple graph of degree three corresponds to the smallest cycle that can exist in an undirected graph. Therefore, if the edges are oriented in the same direction, a cycle of order three can be created among the container adjacent to $v_{0}$. If we look at this situation further, we see that in this example, a node that is an out-neighbor of $v_{0}$ had its out-degree (at least) double. This section will initiate the examination of the adequate conditions for other nodes in the graph to not have their degrees double.
However, we will need to define new concepts to proceed with this line of reasoning. To prevent a node’s out-degree from doubling, we must first formally define the decreasing sequence property. This property is attributed to a node in $G$ whose out-degree does not double in $G^{2}$. There are instances where $v_{0}$ lacks its out-degree double in $G^{2}$, as we have seen. We’ll investigate which other nodes can possess the decreasing sequence property and which network features are necessary for this to occur.